Television sets (particularly older models) produce a ``whistle'' sound at about 15900 Hz (this is the frequency that the electron beam sweeps across the face of the screen). This sound is audible to most youngsters and adults who haven't lost their high-frequency hearing. Imagine that a TV set is on (with the ``mute'' on) in your sister's bedroom as you walk by in the hallway. If your sister's door is ajar, leaving an opening 4 cm wide, and your ear is about 1.5 m from the door as you walk by, for about how many centimeters of your walk will you be able to hear the TV ``whistle'' (if you can at all)?
Answer #1

The speed of this sound is:

v = \lambda f

then:

\lambda = \frac{v}{f}

to hear the whistle, the distance must be:

r= 2L\frac{\lambda}{d} = \frac{2Lv}{fd} =\frac{2(1.5)(343)}{15900(0.04)}

r = 1.62 m